OpenStudy (anonymous):

URGENT! If f (x)=bx+c where b and c are constants, find b and c for f (1)=-3 and f (5)=13

OpenStudy (anonymous):

You there, @clawd ?

OpenStudy (anonymous):

\[ f(1) = b(1)+c = -3\\ f(5) = b(5)+c = 13 \]

OpenStudy (anonymous):

This leads to two equations:\[ b+c = -3\\ 5b+c = 13 \]